# A neural-network based nonlinear non-intrusive reduced basis method with online adaptation for parametrized partial differential equations

### v1

@@ 3.3 Online stage: network adaptation
In the offline stage, we have trained three networks: the dimension reduction network $\Phi$, the solution reconstruction network $U$, and the parameter estimation network $\Theta$. In the online stage, the dimension reduction network is frozen, which means that the reduced basis functions obtained in the offline stage are used directly. We seek to adapt the solution reconstruction model $U(\phi_{1},\ldots,\phi_{r};\boldsymbol{\theta}_{U})$ in an unsupervised manner. Recall that the reduced solution $u_{\rm rb}(\mathbf{x})$ is in the form of Eq. (3.2), and we plug it into the discrete PINNs loss (2.18), yielding our online loss function
${L}_{\rm online}(\boldsymbol{\theta}_{U})=\sum_{i=1}^{m_{1}}\left({\mathcal{N}}[U(\phi_{1}(\mathbf{x}_{i}),\ldots,\phi_{r}(\mathbf{x}_{i});\boldsymbol{\theta}_{U})]\right)^{2}\\ +\lambda\sum_{i=1}^{m_{2}}\left({\mathcal{B}}\left[U(\phi_{1}(\mathbf{x}_{i}),\ldots,\phi_{r}(\mathbf{x}_{i});\boldsymbol{\theta}_{U})\right]\right)^{2}+R(\boldsymbol{\theta}),$ (3.6)
where the sampling points $\{\mathbf{x}_{i}\}_{i=1}^{m_{1}}$ and $\{\mathbf{x}_{i}\}_{i=1}^{m_{2}}$ are generated in the same manner as in standard PINNs, and $R(\boldsymbol{\theta})$ is a regularization term that will be defined later.
Now let $\boldsymbol{\theta}^{*}_{U}$ be the output of the parameter estimation model obtained at the offline stage:
$\boldsymbol{\theta}_{U}^{*}=\Theta(\boldsymbol{\mu}^{*};\boldsymbol{\theta}_{\Theta}),$
and it will be used in the two capacities: for regularization and for initialization. First the regularizer $R(\boldsymbol{\theta})$ is defined as:
$R(\boldsymbol{\theta})=\gamma\|\boldsymbol{\theta}_{U}-\boldsymbol{\theta}_{U}^{*}\|_{2}^{2},$
where $\gamma\geqslant 0$ is the regularization coefficient. Here, we adopt a decaying regularization strategy by making $\gamma$ iteration-dependent, i.e., $\gamma=\gamma_{k}$, where $k$ denotes the epoch number. The regularization weight $\gamma_{k}$ is designed to decrease over time, satisfying $\gamma_{k}\to 0$ as $k\to\infty$. Next, Eq. (3.6) is minimized by using $\boldsymbol{\theta}^{*}_{U}$ as the initial guess. The online training procedure and the network architecture are shown in Fig. 2.
It is important to note that online training typically can be carried out very efficiently, and the reason for this is two-fold. First, as mentioned in Section 3.2, the network $U$ has been deliberately kept small in both depth and width, and so the dimensionality of the parameter space is modest. For example, in Section 4.1, a network with a single hidden layer of 5 neurons is used to reconstruct the solution of the Burgers’ equation. Second, in most cases, $\boldsymbol{\theta}_{U}^{*}$ provides a good initial guess for the training procedure, so that the number of iterations can be significantly reduced.
@@ 4 Numerical experiments
We present two numerical examples in this section, with implementation details provided in the Appendix.
@@ 4.1 Burgers’ equation
$\begin{split}u(x,\boldsymbol{\mu})\frac{\partial u(x,\boldsymbol{\mu})}{\partial x}-\frac{\partial u^{2}(x,\boldsymbol{\mu})}{\partial x^{2}}&=s(x,\boldsymbol{\mu}),\\ u(-1,\boldsymbol{\mu})=u(1,\boldsymbol{\mu})=0,\end{split}$ (4.1)
where $s(\mathbf{x},\boldsymbol{\mu})$ is the source term and $\boldsymbol{\mu}$ is a two dimensional parameter $\boldsymbol{\mu}=[\mu_{1},\mu_{2}]^{T}$ following a uniform distribution defined on $1,10]\times[1,10]$. In this example we choose the source term to be
$s(\mathbf{x},\boldsymbol{\mu})$ $=(1+\mu_{1}x)^{2}\Bigl[\,6x^{2}-2-\frac{2\,h^{2}\mu_{2}^{2}}{9}\,(x^{2}-1)^{2}\Bigr]\cos\!\Bigl(\tfrac{2h\mu_{2}x}{3}\Bigr)$
$\quad+\;\frac{2h\mu_{2}}{3}\,(1+\mu_{1}x)\,(x^{2}-1)\,\bigl[\mu_{1}(x^{2}-1)+2x(1+\mu_{1}x)\bigr]\,\sin\!\Bigl(\tfrac{2h\mu_{2}x}{3}\Bigr)$
@@
In all numerical experiments, we ensure that the set of snapshots is sufficiently large so that it does not become a major limiting factor in performance: the sample size is 100 for $\kappa=1$, and 1600 for $\kappa=9$. We first examine the network model obtained solely from the offline computation. In Fig. 4, we plot the relative mean-squared errors of the reduced solutions versus the number of basis functions for both $\kappa=1$ and $\kappa=9$. The errors are computed using 100 randomly generated test samples. In addition to the results of the proposed method, we show two reference results for comparison. The first is the optimal linear projection, obtained by projecting the exact solution onto the reduced basis; it represents the best-case scenario when a linear mapping is used. The second is the result of the non-intrusive RB method, which, as discussed earlier, is also based on linear mappings. For the less oscillatory case ($\kappa=1$), we observe that the error in the optimal linear projection decays rapidly, becoming smaller than that of our method when six or more basis functions are used. This suggests that in such a simple setting, a linear mapping is capable of approximating the solution quite well. Obviously the optimal projection serves as an ideal benchmark, rather than a practical method. When comparing the results of our nonlinear RB method with those of the linear-mapping-based non-intrusive RB, we can see that the former consistently yields smaller errors, demonstrating the advantage of incorporating nonlinearity in the reduced model. The advantage of the nonlinear mapping becomes more evident in the case of $\kappa=9$. In this more challenging setting, our method not only significantly outperforms the non-intrusive RB method, but also achieves considerably lower errors than the optimal linear projection when using up to 20 basis functions. This highlights the effectiveness of nonlinear representations in capturing complex solution behavior.
@@
Figure 4: Relative errors of the offline-computed solutions as a function of the number of basis functions. Left for the case $\kappa=1$, and right for the case $\kappa=9$.
Next, we consider the effect of online network adaptation, which is a key distinction between our method and the non-intrusive RB method. The first question we aim to address is whether such online computation is necessary. To this end, we plot in Fig. 5 the histogram of the relative errors of the reduced solutions obtained by the offline training only. For $\kappa=1$, we use $r=8$ basis functions, and for $\kappa=9$, we set $r=16$. The results show that the offline model performs very well in the simpler case ($\kappa=1$), but less effectively in the more challenging case ($\kappa=9$). If we use $10^{-3}$ as a threshold for acceptable accuracy, nearly all test cases—94 out of 100—satisfy this criterion for $\kappa=1$, while for $\kappa=9$, only 3 out of 100 cases fall below it. From this we can conclude that online adaptation is particularly important for more complex solutions. Therefore, let us now focus on examining the effect of online adaptation in the $\kappa=9$ case. In Fig. 6 we compare the errors of the reduced solutions obtained by the offline training only and those updated with online adaptation. The left panel shows the average errors for all 100 test samples, while the right panel focuses on the worst cases – that is, the 10% of samples with the largest offline errors. In both cases, online adaptation leads to improved accuracy of the RB solution, with the improvement being particularly significant for the worst cases where the offline model alone performs poorly. These results demonstrate that online adaptation is highly useful, especially when the offline model fails to capture the solution accurately.
@@
Figure 5: Histograms of the relative errors of the offline-computed solutions over 100 test samples. Left for the case $\kappa=1$ using 8 basis functions; right for the case $\kappa=9$ using 16 basis functions.
(a) Online vs Offline: relative errors averaged over all test samples
(b) Online vs Offline: relative errors averaged over 1010 worst cases
Figure 6: Relative errors of the solutions reconstructed using the offline stage only versus those with online adaptation.
4.2 Lid-driven cavity problem
In this section, we consider a more challenging example—the lid-driven cavity problem, which is a widely used benchmark in computational fluid dynamics [1, 2]. We are looking to solve a parametrized version of this problem, which is a challenging task for the following reasons: (1) it involves a complex and computationally intensive parameter-dependent PDE system, (2) the computational domain and the boundary conditions are parameter dependent, and (3) the associated parameter domain is high-dimensional. Specifically, we are looking to compute the velocity and pressure fields of a viscous flow within a parallelogram-shaped domain (where the angle between the horizontal and slanted edges is a random parameter), subject to parametrized boundary conditions. This problem has been used as a test example in several RB-related works such as [18, 8].
The problem is governed by the following equations:
[…]
where $\mathbf{u}=(u,v)^{\top}$ is the velocity field, $p$ is the pressure, $Re$ is the Reynolds number, and $\Omega(\theta)$ denotes the computational domain. The domain $\Omega(\theta)$ is a parallelogram centered at the origin, with the length of all edges being 1, and $\theta$ being the angle between the base and the slanted edge (see Fig. 7). The boundary of the domain, denoted by $\varGamma$, is composed of the lid (i.e., the top horizontal edge) $\varGamma_{2}$ and the walls (the three other edges of the parallelogram) $\varGamma_{1}$. The boundary conditions are given by
[…]
and
$v(\boldsymbol{x})=0,\quad\boldsymbol{x}\in\varGamma,$
@@ 4.2 Lid-driven cavity problem
$R(x)=0.001\sum_{i=1}^{10}\alpha_{i}\sin(i\pi x)\cos(i\pi x).$ (4.4)
The parameters in this PDE model are the domain angle $\theta$, the Reynolds number $Re$, and the boundary condition coefficients $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{10})$; thus, overall, the parameter domain is twelve-dimensional. All these parameters are the images of independent uniformly distributed random variables; specifically, $\theta\sim U[\frac{\pi}{3},\frac{2\pi}{3}]$, $Re\sim U[100,200]$, and $\boldsymbol{\alpha}\sim U[-1,1]^{10}$.
Figure 7: Examples of the computational domain $\Omega(\theta)$ with $\theta\in\{\pi/3,\,\pi/2,\,2\pi/3\}$.
Refer to caption Refer to caption Refer to caption
Figure 8: Relative errors of the offline-computed solutions as a function of the number of basis functions. From left to right, the plots show the results for horizontal velocity, vertical velocity, and pressure.
Figure 9: Histograms of the relative errors (over 100 samples) from the offline training results. The left column corresponds to predictions trained with 100 snapshots, and the right column to those with 300 snapshots. From top to bottom, the rows show the results for the horizontal velocity, vertical velocity, and pressure.
For computationally expensive parametrized PDEs such as the one in this example, it is often impractical to assume that a sufficient number of high-fidelity solutions (i.e., snapshots) can be generated. Therefore, it is informative to evaluate the performance of RB methods under varying number of available snapshots. In our experiments, we consider two snapshot sizes: 100 and 300.
We first consider the offline reconstruction NN model and compare its results with the non-intrusive RB method, using 100 randomly generated realizations of the PDE parameters. In Fig. 8, we plot the relative errors of both methods against the number of basis functions. This figure shows that for all solution components (i.e., for $u$, $v$ and $p$), the approximations generated from 300 snapshots have smaller relative errors than the approximations obtained from 100 snapshots. This indicates that the performance of all tested reduced models is sensitive to the snapshot size. Moreover, we can see from Fig. 8 that the nonlinear RB method consistently achieves lower errors than the non-intrusive RB method for a fixed snapshot size, suggesting that it can capture the underlying solution structure more effectively in this example. Another aspect illustrated in Fig. 8 is how the size of the reduced basis affects the method’s performance: while increasing the size of the reduced basis generally reduces the error, especially for the nonlinear RB method, non-intrusive RB methods can exhibit strong non-monotonic behaviour, particularly when the number of snapshots is limited. These results show that the nonlinear RB method is more robust with respect to the snapshot size, largely due to its more flexible dependence on the snapshot data. Finally, the errors in all cases remain relatively large, which suggests that further online training is necessary to improve the accuracy of the reduced model. This point can be further illustrated by Fig. 9, which shows the histograms of the relative errors for $u$, $v$, and $p$, respectively. If we use $10^{-3}$ as a threshold, the relative errors for a large number of test samples exceed this value, especially in the case with 100 snapshots. These results further confirm that offline computation alone is often insufficient to ensure high accuracy, particularly when the number of snapshots is limited.
Next, we examine the results of the online adaptation using the same 50 test samples that were used to evaluate the offline results. First, as mentioned in Section 3, an important task of the offline model is that it can provide good initialization for the online adaptation. This is particularly important for complex problems, where online training can be highly expensive. We now demonstrate the effectiveness of this strategy through numerical test results. Namely, we conduct online adaptation on three examples that are randomly chosen from the test set and plot the validation loss as a function of the number of epochs in Fig. 10 . The results show how fast the reconstruction models improve during the online training process. In particular, we conclude that the offline prediction provides a significantly lower loss compared to the tested random initialization. More importantly, it leads to much faster convergence: if we use $10^{-4}$ as the stopping threshold, the loss drops below $10^{-4}$ within $10^{4}$ epochs in test examples 2 and 3, and within $4\times 10^{4}$ epochs in example 1, whereas the random initialization converges more slowly and remains above the $10^{-4}$ threshold for all three examples. We also aim to provide a visual comparison of the online and offline results using these three examples. In Fig. 11, we present the velocity field and the pressure for these samples as computed by the offline model, the online model, and the high-fidelity solver. In Fig. 12, we show the spatial error maps in pressure for both the offline and online models, evaluated against the high-fidelity solutions. The plots demonstrate that in the three examples, online adaptation yields clear improvements.
We now evaluate the performance of the online adaption over all the 100 test samples. In Fig. 13, we plot the relative errors of the offline-only predictions and those with online adaptation, against the number of RB functions. The results are shown for the average error over all test samples and the average error over the 10 worst-case samples having the largest offline prediction errors. We present these results for both the 100-snapshot case (Fig. 13(a)) and the 300-snapshot case (Fig. 13(b)). As shown in the figure, in most cases the online adaptation further improves the accuracy of the reduced models. The improvement is particularly evident for the worst-case samples, where offline errors are the largest. Moreover, the effect of online adaptation is clearly more significant in the 100-snapshot case, indicating that online adaptation becomes particularly necessary when snapshot data are limited.
Finally, we discuss the computational cost of solving the parametrized PDE problem in this example. We report wall-clock times based on our specific computational setup, as detailed in Appendix A. Generating a single high-fidelity snapshot typically takes between 4 to 5 hours. The time required for offline training is reported in Table 1. The time for online adaptation varies across different test samples, and Fig. 14 shows the histograms of online training times for the 50 test examples, for both the 100-snapshot (left) and 300-snapshot (right) cases. The figure shows that, in both cases, the online training takes no more than 1600 seconds, which is considerably more efficient than using the conventional solver that typically requires several hours. Moreover, the online training time in the 100-snapshot case is generally longer than in the 300-snapshot case, indicating that when the number of snapshots is limited, online adaptation is particularly needed for achieving reliable results.
$r=5$ $r=10$ $r=15$ $r=20$
300 snapshots 6 15 21 21
100 snapshots 4 7 7 7
Table 1: Offline training time (in hours).
Figure 10: Online training loss plotted against the number of epochs for three test examples.
Figure 11: The velocity-pressure solutions to the lid-driven cavity problem. Each column corresponds to one of three test examples. From top to bottom, three rows show the results from the offline-computed model, the online-computed model, and the high-fidelity reference solution. The streamlines represent the velocity field (flow direction and structure), while the color map represents the pressure distribution.
Figure 12: Relative error maps for the pressure in three test examples: the top row shows the results of the offline only predictions, while the bottom row shows the results where online adaptation is included.
(a) Results for the case with 100 snapshots
(b) Results for the case with 300 snapshots
Figure 13: Relative errors in the solution components reconstructed using the offline stage only versus those with online adaptation.
@@
Figure 14: Histograms of online training times for 50 test examples. Left: the results for 100 snapshots and 10 basis functions. Right: the results for 300 snapshots and 15 basis functions.

### v2

@@ 3.3 Online adaptation
In the offline stage, we have trained three networks: the DR network $\Psi$, the SA network $\Phi$, and the hypernet $\sigma$. In the online stage, the DR network $\Psi$ is frozen, which means that the feature functions obtained in the offline stage are used directly. We seek to adapt the solution reconstruction model $\Phi(\boldsymbol{\psi};\boldsymbol{\theta}_{\mathrm{SA}})$ in an unsupervised manner. With the feature map $\boldsymbol{\psi}$ being fixed, the NNO solution in the online stage can be written as
$u_{\mathrm{on}}(\mathbf{x};\boldsymbol{\theta}_{\mathrm{SA}})=\Phi(\boldsymbol{\psi};\boldsymbol{\theta}_{\mathrm{SA}})(\mathbf{x})=h(\boldsymbol{\psi}(\mathbf{x});\boldsymbol{\theta}_{\mathrm{SA}}),$ (3.7)
and is trained using the PINNs loss (2.15) with a collection of pre-selected test functions. To accelerate the computation, the optimization problem is initialized using the output of the hypernet model obtained at the offline stage:
$\boldsymbol{\theta}_{\mathrm{SA}}^{*}=\sigma(\boldsymbol{\mu}^{*};\boldsymbol{\theta}_{\mathrm{HN}}).$
The online training procedure and the network architecture are shown in Figure 2.
It is important to note that online training typically can be carried out very efficiently, owing to several factors. First, as mentioned in Section 3.2, the network $\Phi$ has been deliberately kept small in both depth and width, and so the dimensionality of the parameter space is modest. For example, in Section 4.1, a network with a single hidden layer of 5 neurons is used to reconstruct the solution of the Burgers’ equation. Second, the evaluation and optimization of the PINN loss requires the computation of the reconstructed NN-based solution and its derivatives (with respect to both spatial variables and network parameters) at a large number of collocation points. Following Proposition 2, the NNO formulation provides pointwise explicit expressions for both the spatial and the parameter gradients, considerably reducing the computational cost of optimizing the loss. Finally, in most cases, $\boldsymbol{\theta}_{\mathrm{SA}}^{*}$ provides a good initial guess for the training procedure, so that the number of iterations can be significantly reduced.
@@
3.4 Connections to PINNs
We briefly discuss the motivation of the NNO method from the PINN perspective. PINNs are often computationally intensive when applied to complex PDE models: firstly, in order to accurately approximate solutions to such models, the neural network typically needs to be large in both depth and width, and secondly, a substantial number of training iterations is often required to achieve a satisfactory solution of the associated minimization problem. In what follows, we will discuss how the proposed method can mitigate this computational burden by introducing an offline pretraining stage, which reduces the complexity of the online PINN training.
The goal of PINNs is to construct a scalar-valued function of the physical variables that approximates the true solution. In this sense, PINNs can be seen as analogous to classical regression problems, where one approximates an underlying target function within a prescribed model class. In regression, feature maps are often introduced to simplify the approximation of complicated and highly nonlinear target functions by representing them in a more expressive or structured feature space. In this regard, the main idea of the NNO method is to construct an explicit feature map and to approximate the solution as a scalar-valued function defined on the resulting feature space, rather than directly on the physical space. The underlying motivation is that, in the feature space, the input–output mapping may admit a simpler representation than in the original physical variables, allowing the solution to be represented using a much smaller network than in standard PINNs. In our method, such effective feature representations are learned from high-fidelity solution snapshots during the offline stage. During the online stage, the solution approximation is then carried out in the resulting feature space using the precomputed feature representation. In summary, an alternative interpretation of the proposed NNO approach is a feature-based PINN approach implemented within an offline–online computational framework.
@@ 4 Numerical experiments
We test the proposed NNO method in two numerical examples, with implementation details provided in the Appendix A.
@@ 4.1 Burgers’ equation
$\begin{split}&u\frac{\partial u}{\partial x}-\frac{\partial^{2}u}{\partial x^{2}}=s(x,\boldsymbol{\mu}),\\ &u(-1)=u(1)=0,\end{split}$ (4.1)
where $s(x,\boldsymbol{\mu})$ is the source term and $\boldsymbol{\mu}$ is a two-dimensional parameter $\boldsymbol{\mu}=[\mu_{1},\mu_{2}]^{T}$ following a uniform distribution defined on $1,10]\times[1,10]$. In this example we choose the source term to be
$s(x,\boldsymbol{\mu})$ $=(1+\mu_{1}x)^{2}\Bigl[\,6x^{2}-2-\frac{2\,\kappa^{2}\mu_{2}^{2}}{9}\,(x^{2}-1)^{2}\Bigr]\cos\!\Bigl(\tfrac{2\kappa\mu_{2}x}{3}\Bigr)$
$\quad+\;\frac{2\kappa\mu_{2}}{3}\,(1+\mu_{1}x)\,(x^{2}-1)\,\bigl[\mu_{1}(x^{2}-1)+2x(1+\mu_{1}x)\bigr]\,\sin\!\Bigl(\tfrac{2\kappa\mu_{2}x}{3}\Bigr)$
@@
In all numerical experiments, we ensure that the set of snapshots is sufficiently large so that it does not become a major limiting factor in performance: the sample size is 100 for $\kappa=1$, and 1600 for $\kappa=9$. We first examine the NNO model obtained solely from the offline computation. In Fig. 4, we plot the relative mean-squared errors of the reduced solutions versus the number of basis functions for both $\kappa=1$ and $\kappa=9$. The errors are computed using 100 randomly generated test samples. In addition to the results of the proposed method, we show two reference results for comparison. The first is the optimal linear projection, obtained by projecting the exact solution onto the reduced basis; it represents the best-case scenario when a linear mapping is used. The second is the result of the non-intrusive RB method, which, as discussed earlier, is also based on linear combination. For the less oscillatory case ($\kappa=1$), we observe that the error in the optimal linear projection decays rapidly, becoming smaller than that of our method when six or more basis functions are used. This suggests that in such a simple setting, a linear combination is capable of approximating the solution quite well. Obviously the optimal projection serves as an ideal benchmark, rather than a practical method. When comparing the results of the NNO method with those of the linear-combination-based non-intrusive RB, we can see that the former consistently yields smaller errors, demonstrating the advantage of incorporating nonlinearity in the reduced model. The advantage of the nonlinear NNO becomes more evident in the case of $\kappa=9$. In this more challenging setting, our method not only significantly outperforms the non-intrusive RB method, but also achieves considerably lower errors than the optimal linear projection when using up to 20 basis/feature functions. This highlights the effectiveness of NNO in capturing complex solution behavior.
@@
Figure 4: Relative errors of the offline-computed solutions as a function of the number of basis/feature functions. Left for the case $\kappa=1$, and right for the case $\kappa=9$.
Next, we consider the effect of online network adaptation. The first question we aim to address is whether such online computation is necessary. To this end, we plot in Fig. 5 the histogram of the relative errors of the reduced solutions obtained by the offline training only. For $\kappa=1$, we use $r=8$ basis/feature functions, and for $\kappa=9$, we set $r=16$. The results show that the offline model performs very well in the simpler case ($\kappa=1$), but less effectively in the more challenging case ($\kappa=9$). If we use $10^{-3}$ as a threshold for acceptable accuracy, nearly all test cases—94 out of 100—satisfy this criterion for $\kappa=1$, while for $\kappa=9$, only 3 out of 100 cases fall below it. From this we can conclude that online adaptation is particularly important for more complex solutions. Therefore, let us now focus on examining the effect of online adaptation in the $\kappa=9$ case. In Fig. 6 we compare the results of the NNO solutions obtained by the offline training only and those updated with online adaptation. Note that in the online adaptation stage, the total number of training iterations is fixed to 100. The left panel shows the average residual loss over 100 test samples before and after online adaptation. The right panel compares the average relative errors of the same set of samples before and after online adaptation. As expected, the left panel shows that online training substantially reduces the residual loss, demonstrating the effectiveness of residual loss–driven unsupervised learning. More importantly, the right panel shows that this reduction translates into a significant improvement in solution accuracy, by approximately one order of magnitude. These results demonstrate that online adaptation is highly effective in improving the accuracy of the NNO solution, particularly for parameter values that are insufficiently represented in the offline training data.
@@
Figure 5: Histograms of the relative errors of the offline-trained model over 100 test samples. Left for the case $\kappa=1$ using 8 feature functions; right for the case $\kappa=9$ using 16 feature functions.
(a) Online vs Offline: residual losses averaged over all test samples
(b) Online vs Offline: relative errors averaged over all test samples
Figure 6: Relative errors and residual losses of the solutions reconstructed using the offline stage only versus those with online adaptation.
4.2 Nonlinear heat conduction problem
We consider a 3-dimensional steady–state nonlinear heat conduction problem
$\nabla\cdot\bigl(k(u)\nabla u\bigr)+S(\mathbf{x};\boldsymbol{\mu})=0,\qquad\mathbf{x}=(x_{1},x_{2},x_{3})\in\Omega,$ (4.3)
posed on the cubic domain
$\Omega=(0,L)^{3}\subset\mathbb{R}^{3},$
with homogeneous Dirichlet boundary conditions
$u=0\qquad\text{on }\partial\Omega.$ (4.4)
Here, $u(\mathbf{x})$ denotes the temperature field, and the thermal conductivity is temperature-dependent and given by
$k(u)=k_{0}(1+\beta u),$ (4.5)
where $k_{0}>0$ and $\beta\geq 0$ are constants. Here $k_{0}>0$ is the reference conductivity and $\beta\geq 0$ controls the strength of the nonlinearity. When $\beta=0$, the problem reduces to a linear Poisson equation, whereas $\beta>0$ introduces a nonlinear diffusion effect. The volumetric heat source is spatially varying but independent of the temperature:
$S(\mathbf{x};\boldsymbol{\mu})=\sum_{i=1}^{N_{S}}\mu_{i}\sin\!\left(\frac{i\pi x_{1}}{L}\right)\sin\!\left(\frac{i\pi x_{2}}{L}\right)\sin\!\left(\frac{i\pi x_{3}}{L}\right),$ (4.6)
@@ 4.2 Nonlinear heat conduction problem
$\boldsymbol{\mu}=(\mu_{1},\dots,\mu_{N_{S}}),$
is a vector of random parameters. In this example, we take $k_{0}=1$, $\beta=2$, $L=2$, $N_{S}=12$ and $\boldsymbol{\mu}$ to follow a uniform distribution $U[-1,1]^{N_{S}}$. High-fidelity FEM snapshots are obtained using eight-node linear hexahedral elements and a Newton-Raphson solver, with the numerical configuration tuned to ensure the relative error is controlled below $10^{-6}$. We generate 1000 snapshots for the offline training, for both the basis generation and network training.
@@
Figure 7: Relative errors of the offline-trained model as a function of the number of basis/feature functions. Left: Errors averaged across all 100 test samples; Right: Errors averaged across the 10 worst cases.
Refer to caption Refer to caption
Figure 8: Relative errors of the solutions reconstructed using the offline stage only versus those with online adaptation. Left: Errors averaged across all 100 test samples; Right: Errors averaged across the 10 worst cases.
We first consider the offline NNO model and compare its results with the non-intrusive RB method, using 100 randomly generated realizations of the PDE parameters. In Fig. 7 (left), we present the relative errors of the NNO method and the non-intrusive RB method as functions of the number of basis functions, averaged over 100 randomly generated test samples. The results show that the nonlinear RB method consistently outperforms the non-intrusive RB approach across all tested basis sizes, indicating that the nonlinear construction is more effective at capturing the intrinsic structure of the solution manifold in this example. Figure 7 (right) further reports the worst 10 cases among the 100 test samples. Although the overall trends remain similar, the errors in these challenging instances are substantially larger than the averaged results. In particular, even for relatively large basis dimensions, the worst-case error of NNO remains on the order of $10^{-2}$, while that of the non-intrusive RB method is considerably higher. This observation indicates that the offline-constructed NN models are not sufficiently robust to uniformly approximate all parameter realizations.
Next, we examine the results of the online adaptation using the same 100 test samples that were used to evaluate the offline results. In Fig. 8, we plot the relative errors of the offline-only predictions and those with online adaptation, against the number of RB functions. The results are shown for the average error over all test samples (left plot) and the average error over the 10 worst-case samples having the largest offline prediction errors (right plot). As shown in the figures, the online adaptation substantially improves the accuracy of the NNO models. The improvement is particularly significant for the worst-case samples, where the offline errors are largest. In these cases, the average error is reduced from above $1.2\times 10^{-2}$ to less than $4\times 10^{-3}$, bringing it well below the $10^{-2}$ level.
Finally, we discuss the computational cost of solving the parametrized PDE problem in this example. We report wall-clock times based on our specific network architecture, as detailed in Appendix A. Generating a single high-fidelity snapshot typically takes approximately two hours. The offline training phase requires about 20 hours for configurations with both 12 and 18 feature functions. The time for online adaptation varies across different test samples, and Fig. 9 shows the histograms of online training times for the 100 test examples, for two different numbers of feature functions 12 and 18. The figure shows that, for both configurations, the online training takes less than 50 seconds, which is considerably more efficient than using the conventional solver. We note that, although offline training takes significantly longer than online training, the primary offline computational cost comes from high-fidelity snapshot generation, similar to non-intrusive RB methods.
Figure 9: Histograms of online training times for 100 test examples. Left: the results for 12 feature functions. Right: the results for 18 feature functions.